# Transcoders Find Interpretable LLM Feature Circuits

### v1

@@ Abstract
A key goal in mechanistic interpretability is circuit analysis: finding sparse subgraphs of models corresponding to specific behaviors or capabilities. However, MLP sublayers make fine-grained circuit analysis on transformer-based language models difficult. In particular, interpretable features—such as those found by sparse autoencoders (SAEs)—are typically linear combinations of extremely many neurons, each with its own nonlinearity to account for. Circuit analysis in this setting thus either yields intractably large circuits or fails to disentangle local and global behavior. To address this we explore transcoders, which seek to faithfully approximate a densely activating MLP layer with a wider, sparsely-activating MLP layer. We successfully train transcoders on language models with 120M, 410M, and 1.4B parameters, and find them to perform at least on par with SAEs in terms of sparsity, faithfulness, and human-interpretability. We then introduce a novel method for using transcoders to perform weights-based circuit analysis through MLP sublayers. The resulting circuits neatly factorize into input-dependent and input-invariant terms. Finally, we apply transcoders to reverse-engineer unknown circuits in the model, and we obtain novel insights regarding the “greater-than circuit” in GPT2-small. Our results suggest that transcoders can prove effective in decomposing model computations involving MLPs into interpretable circuits. Code is available at https://github.com/jacobdunefsky/transcoder_circuits.
@@ 1 Introduction
A core problem in fine-grained circuit analysis is incorporating MLP sublayers [12, 10]. Attempting to analyze MLP neurons directly suffers from “polysemanticity” [13, 14, 15, 16]: the tendency of neurons to activate on many unrelated concepts. To address this, sparse autoencoders (SAEs) [17, 18, 19] have been used to perform fine-grained circuit analysis by instead looking at features—vectors in the model’s representation space—instead of individual neurons [20, 21]. However, while SAE features are often interpretable, these vectors tend to be dense linear combinations of many neurons [22]. Thus, mechanistically understanding how an SAE feature before one or more MLP layers affects a later SAE feature may require considering an infeasible number of neurons and their nonlinearities. Prior attempts to circumvent this [20, 21] use a mix of causal interventions and gradient-based approximations to MLP layers. But these approaches fail to exhibit input-invariance: the connections between features can only ever be described for a given input, and not for the model as a whole. Attempts to address this, e.g. by averaging results over many inputs, conversely lose their ability to yield input-dependent information that describes a connection’s importance on a single input.
@@
Motivated by this, in this work, we explore transcoders (an idea proposed, but not explored, in Templeton et al. [23] and Li et al. [24]): wide, sparsely-activating approximations of a model’s original MLP sublayer. Specifically, an MLP transcoder is a wide ReLU MLP sublayer (with one hidden layer) that is trained to faithfully approximate the original narrower MLP sublayer’s output, with an L1 regularization penalty on neuron activations to encourage sparse activations. The main advantage of transcoders is they replace a difficult-to-interpret model component—the MLP sublayer—with an interpretable approximation that is otherwise faithful to the original computations. This allows us to interpret transcoder neurons rather than dense linear combinations of original MLP neurons.
Our contributions.
Our main contributions are (1) to confirm that transcoders are a faithful and interpretable approximation to MLP sublayers, and (2) to demonstrate a novel method for circuit analysis using transcoders.
In §3.2, we evaluate transcoders’ interpretability, sparsity, and faithfulness to the original model. Because SAEs are the standard method for finding sparse decompositions of model activations, we compare transcoders to SAEs on models up to 1.4 billion parameters and verify that transcoders are on par with SAEs or better with respect to these properties.
Beyond this, however, transcoders additionally enable circuit-finding techniques that are not possible using SAEs: in §4.1 we introduce a novel method for performing circuit analysis with transcoders and demonstrate that transcoders cleanly factorize circuits into input-invariant and input-dependent components. We apply transcoder circuit analysis to a variety of tasks in §4.2 and §4.3, including “blind case studies,” which demonstrate how this approach allows us to understand features without looking at specific examples, and an in-depth analysis of the GPT2-small “greater-than circuit” previously studied by Hanna et al. [25].
Code for training transcoders and carrying out our experiments is available at https://github.com/jacobdunefsky/transcoder_circuits.
@@ 2 Transformers preliminaries
Following the approach of Elhage et al. [9], the computation of a transformer model can be represented as follows. First, the model maps input tokens (and their positions) to embeddings $\mathbf{x_{pre}^{(0,t)}}\in\mathbb{R}^{d_{\text{model}}}$, where $t$ is the token index and $d_{\text{model}}$ is the model dimensionality. Then, the model applies a series of “layers,” which map the hidden state at the end of the previous block to the new hidden state. This can be expressed as:
@@
where $l$ is the layer index, $t$ is the token index, $\operatorname{attn}^{(l,h)}(\mathbf{x^{(l,t)}_{pre}};\mathbf{x^{(l,1:t)}_{pre}})$ denotes the output of attention head $h$ at layer $l$ given all preceding source tokens $\mathbf{x^{(l,1:t)}_{pre}}$ and destination token $\mathbf{x^{(l,t)}_{pre}}$, and $\operatorname{MLP}^{(l)}(\mathbf{x^{(l,t)}_{mid}})$ denotes the output of the layer $l$ MLP.
@@ 3.1 Architecture and training
Transcoders aim to learn a “sparsified” approximation of an MLP sublayer: a transcoder approximates the output of an MLP sublayer as a sparse linear combination of feature vectors. Formally, the transcoder architecture can be expressed as
@@
Each feature in a transcoder is associated with two vectors: the $i$-th row of $\mathbf{W_{enc}}$ is the encoder feature vector of feature $i$, and the $i$-th column of $\mathbf{W_{dec}}$ is the decoder feature vector of feature $i$. The $i$-th component of $\mathbf{z_{TC}(\mathbf{x})}$ is called the activation of feature $i$. Intuitively, for each feature, the encoder vector is used to determine how much the feature should activate; the decoder vector is then scaled by this amount, and the resulting weighted sum of decoder vectors is the output of the transcoder. In this paper, the notation $\mathbf{f^{(l,i)}_{enc}}$ and $\mathbf{f^{(l,i)}_{dec}}$ are used to denote the $i$-th encoder feature vector and decoder feature vector, respectively, in the layer $l$ transcoder.
Because we want transcoders to learn to approximate an MLP sublayer’s computation with a sparse linear combination of feature vectors, transcoders are trained with the following loss, where $\lambda_{1}$ is a hyperparmeter mediating the tradeoff between sparsity and faithfulness:
@@
3.1.1 Evaluation metrics
We evaulate transcoders qualitatively on their features’ interpretability as judged by a human rater, and quantitatively according to the sparsity of their activations and their fidelity to the original MLP’s computation.
As a qualitative proxy measure for the interpretability of a feature, we follow Bricken et al. [17] in assuming that interpretable features should demonstrate interpretable patterns in the examples that cause them to activate. To this end, one can run the transcoder on a large dataset of text, see which dataset examples cause the feature to activate, and see if there is an interpretable pattern among these tokens. While this is an imperfect metric [26], it is still a reasonable proxy for an inherently qualitative concept.
3.2 Relationship to SAEs
Transcoders were originally conceived as a variant of SAEs, and as such, there are many similarities between them. SAEs have the exact same architecture as transcoders in the sense that they also have encoder and decoder feature vectors, but differ in how they are trained: because SAEs are autoencoders, the faithfulness term in the SAE loss measures the reconstruction error between the SAE’s output and its original input. In contrast, the faithfulness term of the transcoder loss measures the error between the transcoder’s output and the original MLP sublayer’s output.
Because of the extensive similarities between SAEs and transcoders, SAEs can be quantitatively evaluated (for sparsity and fidelity) and qualitatively evaluated (for feature interpretability) in precisely the same way as transcoders. In fact, the aforementioned transcoder evaluation methods are also standard for evaluating SAEs [27, 28]. We now report the results of evaluations comparing SAEs to transcoders on these metrics, and find that transcoders are comparable to or better than SAEs.
3.2.1 Blind interpretability comparison of transcoders to SAEs
In order to evaluate the interpretability of transcoders, we manually attempted to interpret 50 random features from a Pythia-410M [29] layer 15 transcoder and 50 random features from a Pythia-410M layer 15 SAE trained on MLP inputs. For each feature, the examples in a subset of the OpenWebText corpus [30] that caused the feature to activate the most were computed ahead of time. Then, the features from both the SAE and the transcoder were randomly shuffled. For each feature, the maximum-activating examples were displayed, but not whether the feature came from an SAE or transcoder. We recorded for each feature whether or not there seemed to be an interpretable pattern, and only after examining every feature did we look at which features came from where. The results, shown in Table 1, suggest transcoder features are approximately as interpretable as SAE features. This further suggests that transcoders incur no penalties compared to SAEs.
Table 1: The number of interpretable features, possibly-interpretable features, and uninterpretable features for the transcoder and SAE. Of the interepretable features, we additionally deemed 6 transcoder features and 16 SAE features to be “context-free”, meaning they appeared to fire on a single token without any evident context-depdendent patterns.
# interpretable # maybe # uninterpretable
Transcoder 41 8 1
SAE 38 8 4
3.2.2 Quantitative comparison of transcoders to sparse autoencoders
We now compare transcoders to SAEs according to the sparsity and fidelity metrics discussed in §3.1.1. We trained transcoders on MLP activations and SAEs on MLP output activations from GPT2-small [31], Pythia-410M, and Pythia-1.4B. For each model, we trained multiple SAEs and transcoders on the same activations, but varying the $\lambda_{1}$ hyperparameter controlling the fidelity-sparsity tradeoff. We evaluated each SAE and transcoder on the same 3.2 million tokens of OpenWebText data. We also recorded the unmodified model’s loss and the loss after mean-ablating the entire MLP sublayer (always replacing its output with its mean output over the dataset) as best- and worst-case bounds, respectively.
4 Circuit analysis with transcoders
4.1 Circuit analysis method
@@
We now introduce a novel method for performing feature-level circuit analysis with transcoders, which provides a scalable and interpretable way to identify which transcoder features in different layers connect to compute a given task. Moreover, this method neatly factorizes the importance of computational subgraphs into input-invariant terms, which can be computed just from model and transcoder weights, and input-dependent terms, which depend on the specific model input.
4.1.1 Attribution between feature pairs
The primary goal of circuit analysis is to identify a subgraph of the model’s computational graph that is responsible for (most of) the model’s behavior on a given task [32, 33, 34]; this requires a means of evaluating a computational subgraph’s importance to the task in question. Unfortunately, as discussed in §1, MLP sublayers make this difficult. But with more interpretable and sufficiently faithful transcoders, we can replace the MLP sublayers to obtain a more interpretable computational graph more amenable to circuit analysis.
@@ 4.1.1 Attribution between feature pairs
In order to identify the relevant subgraph in this transcoder computational graph, we begin with the following insight: every transcoder feature contributes some (possibly zero) amount to the residual stream, which results in some contribution to all subsequent transcoder features. This means that we can quantify the attribution of an earlier-layer feature to a later-layer feature’s activation—which allows us to identify important edges in the computational graph. This attribution is given by the product of two terms: the earlier feature’s activation (which depends on the input to the model), and the dot product of the earlier feature’s decoder vector with the later feature’s encoder vector (which is independent of the model input).
@@ 4.1.2 Finding computational subgraphs
Doing so iteratively (and greedily pruning all but the most important features at each step) thus yields a set of computational paths (a sequence of connected edges). These computational paths can then be combined into a computational subgraph, in such a way that each node (transcoder feature), edge, and path is assigned an attribution. This process can be further extended to take into account the “OV circuits” of attention heads (under the formalism presented by Elhage et al. [9]) as described in App. D.3. This allows contributions from previous tokens in the input to the current token to be accounted for. A full description of the circuit-finding algorithm is presented in App. D.5.
attn0[5]@37embed0@37attn0[1]@37attn0[3]@37tc0[9188]@37tc0[16632]@371.1tc2[3900]@37tc1[22184]@371.1tc3[6238]@37tc8[355]@372.4
Figure 3: An example of a computational graph produced using the method in §4.1.2 characterizing how our unknown feature is computed on an unseen input. A single path is highlighted in red and annotated with component-by-component attributions.
@@ 4.1.3 De-embeddings: a special case of input-invariant information
Given a de-embedding vector, looking at which tokens in the model’s vocabulary have the highest de-embedding scores tells us about the feature’s general behavior. For example, for a certain GPT2-small MLP0 transcoder feature that we investigated, the tokens with the highest scores were oglu, owsky, zyk, chenko, and kowski. Notice the interpretable pattern: all of these tokens come from European surnames, primarily Polish ones. This suggests that the general behavior of the feature is to fire on Polish surnames.

### v2

@@ Abstract
A key goal in mechanistic interpretability is circuit analysis: finding sparse subgraphs of models corresponding to specific behaviors or capabilities. However, MLP sublayers make fine-grained circuit analysis on transformer-based language models difficult. In particular, interpretable features—such as those found by sparse autoencoders (SAEs)—are typically linear combinations of extremely many neurons, each with its own nonlinearity to account for. Circuit analysis in this setting thus either yields intractably large circuits or fails to disentangle local and global behavior. To address this we explore transcoders, which seek to faithfully approximate a densely activating MLP layer with a wider, sparsely-activating MLP layer. We introduce a novel method for using transcoders to perform weights-based circuit analysis through MLP sublayers. The resulting circuits neatly factorize into input-dependent and input-invariant terms. We then successfully train transcoders on language models with 120M, 410M, and 1.4B parameters, and find them to perform at least on par with SAEs in terms of sparsity, faithfulness, and human-interpretability. Finally, we apply transcoders to reverse-engineer unknown circuits in the model, and we obtain novel insights regarding the “greater-than circuit” in GPT2-small. Our results suggest that transcoders can prove effective in decomposing model computations involving MLPs into interpretable circuits. Code is available at https://github.com/jacobdunefsky/transcoder_circuits/.
@@ 1 Introduction
A core problem in fine-grained circuit analysis is incorporating MLP sublayers [38, 32]. Attempting to analyze MLP neurons directly suffers from “polysemanticity” [40, 16, 3, 24]: the tendency of neurons to activate on many unrelated concepts. To address this, sparse autoencoders (SAEs) [7, 12, 51] have been used to perform fine-grained circuit analysis by instead looking at features—vectors in the model’s representation space—instead of individual neurons [34, 14]. However, while SAE features are often interpretable, these vectors tend to be dense linear combinations of many neurons [36]. Thus, mechanistically understanding how an SAE feature before one or more MLP layers affects a later SAE feature may require considering an infeasible number of neurons and their nonlinearities. Prior attempts to circumvent this [34, 14] use a mix of causal interventions and gradient-based approximations to MLP layers. But these approaches fail to exhibit input-invariance: the connections between features can only ever be described for a given input, and not for the model as a whole. Attempts to address this, e.g. by averaging results over many inputs, conversely lose their ability to yield input-dependent information that describes a connection’s importance on a single input. This means that SAEs cannot tell us about the general input-output behavior of an MLP across all inputs.
To address why input-invariance is desirable, consider the following example: say that one has a post-MLP SAE feature and wants to see how it is computed from pre-MLP SAE features. Doing e.g. patching on one input shows that a pre-MLP feature for Polish last names is important for causing the post-MLP feature to activate. But on other inputs, would features other than the Polish last name feature also cause the post-MLP feature to fire (e.g. an English last names feature)? Could there be other inputs where the Polish last names feature fires but the post-MLP feature does not? We can see that without input-invariance, it is difficult to make general claims about model behavior.
@@
Motivated by this, in this work, we explore transcoders (an idea proposed, but not explored, in Templeton et al. [47] and Li et al. [31]): wide, sparsely-activating approximations of a model’s original MLP sublayer. Specifically, MLP transcoders are wide ReLU MLPs with one hidden layer that are trained to faithfully approximate the original narrower MLP sublayer’s output, with an L1 regularization penalty on neuron activations to encourage sparse activations. Our primary motivation is to enable input-invariant feature-level circuit analysis through MLP sublayers, which allows us to understand and interpret the general behavior of circuits involving MLP sublayers.
@@
Our contributions.
Our main contributions are (1) to introduce a method for circuit analysis using transcoders, (2) to confirm that transcoders are a faithful and interpretable approximation to MLP sublayers, and (3) to demonstrate the utility of our circuit analysis method on detailed case studies.
After describing the architecture of transcoders in §3.1, we demonstrate in §3.2 that transcoders additionally enable circuit-finding techniques that are not possible using SAEs, and introduce a novel method for performing circuit analysis with transcoders and demonstrate that transcoders cleanly factorize circuits into input-invariant and input-dependent components.
Then, in §4, we evaluate transcoders’ interpretability, sparsity, and faithfulness to the original model. Because SAEs are the standard method for finding sparse decompositions of model activations, we compare transcoders to SAEs on models up to 1.4 billion parameters and verify that transcoders are on par with SAEs or better with respect to these properties.
We apply transcoder circuit analysis to a variety of tasks in §5.1 and §5.2, including “blind case studies,” which demonstrate how this approach allows us to understand features without looking at specific examples, and an in-depth analysis of the GPT2-small “greater-than circuit” previously studied by Hanna et al. [26].
@@ 2 Transformers preliminaries
Following Elhage et al. [15], we represent the computation of a transformer model as follows. First, the model maps input tokens (and their positions) to embeddings $\mathbf{x_{pre}^{(0,t)}}\in\mathbb{R}^{d_{\text{model}}}$, where $t$ is the token index and $d_{\text{model}}$ is the model dimensionality. Then, the model applies a series of “layers,” which map the hidden state at the end of the previous block to the new hidden state. This can be expressed as:
@@
where $l$ is the layer index, $t$ is the token index, $\operatorname{attn}^{(l,h)}(\mathbf{x^{(l,t)}_{pre}};\mathbf{x^{(l,1:t)}_{pre}})$ denotes the output of attention head $h$ at layer $l$ given all preceding source tokens $\mathbf{x^{(l,1:t)}_{pre}}$ and destination token $\mathbf{x^{(l,t)}_{pre}}$, and $\operatorname{MLP}^{(l)}(\mathbf{x^{(l,t)}_{mid}})$ denotes the output of the layer $l$ MLP. [^1] Note that the “Pythia” family of models computes MLP and attention sublayer outputs in parallel. This means that Equation 2 is thus given by $\mathbf{x_{pre}^{(l+1,t)}}=\mathbf{x_{mid}^{(l,t)}}+\operatorname{MLP}^{(l)}\left(\mathbf{x^{(l,t)}_{pre}}\right).$
@@ 3.1 Architecture and training
Transcoders aim to learn a “sparsified” approximation of an MLP sublayer: they approximate the output of an MLP sublayer as a sparse linear combination of feature vectors. Formally, the transcoder architecture can be expressed as
@@
Each feature in a transcoder is associated with two vectors: the $i$-th row of $\mathbf{W_{enc}}$ is the encoder feature vector of feature $i$, and the $i$-th column of $\mathbf{W_{dec}}$ is the decoder feature vector of feature $i$. The $i$-th component of $\mathbf{z_{TC}(\mathbf{x})}$ is called the activation of feature $i$. Intuitively, for each feature, the encoder vector is used to determine how much the feature should activate; the decoder vector is then scaled by this amount, and the resulting weighted sum of decoder vectors is the output of the transcoder. In this paper, the notation $\mathbf{f^{(l,i)}_{enc}}$ and $\mathbf{f^{(l,i)}_{dec}}$ is used to denote the $i$-th encoder feature vector and decoder feature vector, respectively, in the layer $l$ transcoder.
Because we want transcoders to learn to approximate an MLP sublayer’s computation with a sparse linear combination of feature vectors, transcoders are trained with the following loss, where $\lambda_{1}$ is a hyperparameter mediating the tradeoff between sparsity and faithfulness:
@@
3.2 Circuit analysis with transcoders
@@
We now introduce a novel method for performing feature-level circuit analysis with transcoders, which provides a scalable and interpretable way to identify which transcoder features in different layers connect to compute a given task. Importantly, this method provides insights into the general input-output behavior of MLP sublayers, which SAE-based methods cannot do.
In particular, the primary goal of circuit analysis is to identify a subgraph of the model’s computational graph that is responsible for (most of) the model’s behavior on a given task [20, 11, 19]; this requires a means of evaluating a computational subgraph’s importance to the task in question. In order to determine which edges are included in this subgraph, we thus have to compute attributions for each edge: how much the earlier node contributes to the later node’s own contribution. Circuit analysis with SAEs thus entails computing the attribution of pre-MLP SAE features to post-MLP SAE features, as mediated through the MLP. Standard methods for computing attributions are causal patching (which inherently only gives information about local MLP behavior on a single input) and methods like input-times-gradient or attribution patching (which are equivalent in this setting). We will now demonstrate why these methods cannot yield information about the MLP’s general behavior. Letting $\mathbf{z}$ be the activation of an earlier-layer feature, $\mathbf{z^{\prime}}$ be the activation of the later-layer feature, and $\mathbf{y}$ be the activation of the MLP at layer $l^{\prime}$, the input-times-gradient is given by:
$\mathbf{z}\left(\frac{\partial\mathbf{z^{\prime}}}{\partial\mathbf{z}}\right)=\mathbf{z}\left(\frac{\partial\mathbf{z^{\prime}}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\mathbf{z}}\right).$ (6)
Unfortunately, not only is $\mathbf{z}$ input-dependent, but so is $\frac{\partial\mathbf{z^{\prime}}}{\partial\mathbf{z}}$ as well, because $\frac{\partial\mathbf{y}}{\partial\mathbf{z}}$ is.
This means that we cannot use SAEs to understand the general behavior of MLPs on various inputs. In contrast, we will show that when we replace MLP sublayers with sufficiently faithful and interpretable transcoders, we obtain attributions that neatly factorize into input-dependent terms and input-invariant terms; the latter can be computed just from model and transcoder weights, and tell us about the MLP behavior across all inputs.
Feature
Step 1: find the direct contributions of earlier-layer features to a later-layer feature.
Feature
Step 2: delete all but the $k$ most highly-contributing features.
FeatureFeature
Step 3: select the top-contributing features for each selected features.
FeatureFeature
Step 4: delete all but the $k$ most important paths in the graph again.
FeatureFeature
Step 5: Repeat steps 2–4 until you have paths of length $l$.
Feature
Step 6: Merge top $k$ paths into a single circuit.
Figure 2: A visualization of the circuit-finding algorithm.
3.2.1 Attribution between transcoder feature pairs
@@
We begin by showing how to compute attributions between pairs of transcoder features. This attribution is given by the product of two terms: the earlier feature’s activation (which depends on the input to the model), and the dot product of the earlier feature’s decoder vector with the later feature’s encoder vector (which is independent of the model input).
@@
3.2.2 Attribution through attention heads
So far, we have addressed how to find the attribution of a lower-layer transcoder feature directly on a higher-layer transcoder feature at the same token. But transcoder features can also be mediated by attention heads. We will thus extend the above analysis to account for finding the attribution of transcoder features through the OV circuit of an attention head. For a full derivation, see App. D.3.
As before, we want to understand what causes feature $i^{\prime}$ in the layer $l^{\prime}$ transcoder to activate on token $t$. Given attention head $h$ at layer $l$ with $l<l^{\prime}$, the contribution of token $s$ at layer $l$ through head $h$ to feature $i^{\prime}$ in layer $l^{\prime}$ at token $t$ is given by
$\operatorname{score}^{(l,h)}\left(\mathbf{x^{(l,t)}_{pre}},\mathbf{x^{(l,s)}_{pre}}\right)\left(\left(\left(\mathbf{W_{OV}^{(l,h)}}\right)^{T}\mathbf{f^{(l^{\prime},i^{\prime})}_{enc}}\right)\cdot\mathbf{x^{(l,s)}_{pre}}\right),$ (8)
where $\operatorname{score}^{(l,h)}\left(\mathbf{x^{(l,t)}_{pre}},\mathbf{x^{(l,s)}_{pre}}\right)$ is the attention score for head $h$ and layer $l$ from token $s$ to token $t$.
@@ 3.2.3 Finding computational subgraphs
Doing so iteratively (and greedily pruning all but the most important features at each step) thus yields a set of computational paths (a sequence of connected edges). These computational paths can then be combined into a computational subgraph, in such a way that each node (transcoder feature or attention head), edge, and path is assigned an attribution. A full description of the circuit-finding algorithm is presented in App. D.5. Figure 2 provides a visualization of this algorithm.
@@ 3.2.4 De-embeddings: a special case of input-invariant information
Given a de-embedding vector, looking at which tokens in the model’s vocabulary have the highest de-embedding scores tells us about the feature’s general behavior. For example, for a certain GPT2-small MLP0 transcoder feature that we investigated, the tokens with the highest scores were oglu, owsky, zyk, chenko, and kowski. Notice the interpretable pattern: all of these tokens come from European surnames, primarily Polish ones, suggesting that the feature generally fires on Polish surnames.
4 Comparison with SAEs
Transcoders were originally conceived as a variant of SAEs, and as such, there are many similarities between them. They differ only in their training objective: because SAEs are autoencoders, the faithfulness term in the SAE loss measures the reconstruction error between the SAE’s output and its original input. In contrast, the faithfulness term of the transcoder loss measures the error between the transcoder’s output and the original MLP sublayer’s output.
Because of these similarities, SAEs can be quantitatively evaluated (for sparsity and fidelity) and qualitatively evaluated (for feature interpretability) in precisely the same way as transcoders, using standard SAE evaluation methods  [29, 4]. We now report the results of evaluations comparing SAEs to transcoders on these metrics, and find that transcoders are comparable to or better than SAEs.
4.1 Blind interpretability comparison of transcoders to SAEs
In order to evaluate the interpretability of transcoders, we manually attempted to interpret 50 random features from a Pythia-410M layer 15 transcoder and 50 random features from a Pythia-410M layer 15 SAE trained on MLP inputs. [^2] We used SAEs trained on MLP inputs here because the interpretability case studies look at feature activations, which are solely dependent on the encoder vectors of the SAEs and transcoders. Because transcoders’ encoder vectors live in MLP input space, we thought that the comparison would be most accurate if our SAEs’ encoder vectors also lived in MLP input space. For each feature, the examples in a subset of the OpenWebText corpus that caused the feature to activate the most were computed ahead of time. Then, the features from both the SAE and the transcoder were randomly shuffled. For each feature, the maximum-activating examples were displayed, but not whether the feature came from an SAE or transcoder. We recorded for each feature whether or not there seemed to be an interpretable pattern, and only after examining every feature did we look at which features came from where. The results, shown in Table 1, suggest transcoder features are approximately as interpretable as SAE features. This further suggests that transcoders incur no penalties compared to SAEs.
Table 1: The number of interpretable features, possibly-interpretable features, and uninterpretable features for the transcoder and MLP-in SAE. Of the interpretable features, we additionally deemed 6 transcoder features, and 16 SAE features to be “context-free”, meaning they appeared to fire on a single token without any evident context-dependent patterns.
Transcoder MLP-in SAE
# interpretable 41 38
# maybe 8 8
# uninterpretable 1 4
4.2 Quantitative comparison of transcoders to SAEs
4.2.1 Evaluation metrics
We evaulate transcoders qualitatively on their features’ interpretability as judged by a human rater, and quantitatively on the sparsity of their activations and their fidelity to the original MLP’s computation.
As a qualitative proxy measure for the interpretability of a feature, we follow Bricken et al. [7] in assuming that interpretable features should demonstrate interpretable patterns in the examples that cause them to activate. To this end, one can run the transcoder on a large dataset of text, see which dataset examples cause the feature to activate, and see if there is an interpretable pattern among these tokens. While imperfect [6], this is still a reasonable proxy for an inherently qualitative concept.
4.2.2 Results
We trained SAEs and transcoders on activations from GPT2-small [44], Pythia-410M, and Pythia-1.4B [2]. For each model, we trained multiple SAEs and transcoders on the same inputs, but with different values of the $\lambda_{1}$ hyperparameter controlling the fidelity-sparsity tradeoff for each SAE and each transcoder. The transcoders were trained on MLP-in and MLP-out activations, while SAEs were trained on MLP-out activations (as these are the activations that MLP SAEs are typically trained on). Due to compute limitations, we used the same learning rate, which was determined via a hyperparameter sweep on transcoders, for both SAEs and transcoders. This means that the learning rate might not be optimal for SAEs. Nevertheless, we did perform a separate hyperparameter sweep of $\lambda_{1}$ for the SAEs and transcoders.
We evaluated each SAE and transcoder on the same 3.2M tokens of OpenWebText data [21]. We also recorded the loss of the unmodified and mean-ablated model (always replacing the MLP sublayer output with its mean output over the dataset) as best- and worst-case bounds, respectively.
5 Circuit analysis case studies